Walking 3/4 of his usual speed , a student is 1o minutes late to his school. find his usual time to cover the distance.
Answer by lwsshak3(11628)

Walking 3/4 of his usual speed , a student is 1o minutes late to his school. find his usual time to cover the distance.
**
let x=usual travel time to cover the distance
let s=usual speed to cover the distance
distance=speed*travel time=sx
distance/speed=travel time
sx/(3/4)s=x+10
4x/3=x+10
x/3=10
x=30
ans:
usual travel time to cover the distance=30 min